Md=25y-80r,Ms=2400, p=2





a)       Find equilibrium output equation. And Y* when r=12%





b)     Find LM curve equation





c)      Find slope of LM curve.





d)     MS increased to 3000





3. Find equilibrium r





4. Find equilibrium y




1
Expert's answer
2022-04-04T09:32:16-0400

Solution

Md=Ms

"2400=25Y-80r"


"r=\frac{25Y-2400}{80}" The LM equation

"25Y=2400-80r"

r=12%


"Y=\frac{2400-9.6}{25}= 95.6"


When Ms is increased to 3000

"3000=25y-80r"

"80= 25y -3000"

LM equation

"\frac{25}{80}-\frac{3000}{80}"


Equilibrium when r=12

"25Y = 3000-80(0.12)"

"Y = 119.6"


c)